What mass of copper(II) hydroxide would be produced when a solution containing 6.8g of OH- ions reacts with a solution containing excess Cu2+ ions?



1
Expert's answer
2021-07-14T05:05:24-0400

Solution:

copper(II) hydroxide = Cu(OH)2

Cu2+ + 2OH- → Cu(OH)2

According to the equation above: n(Cu2+) = n(OH-)/2 = n(Cu(OH)2)


Calculate the moles of OH- ions:

(6.8 g OH-) × (1 mol OH- / 17 g OH-) = 0.4 mol OH-


Thus,

n(Cu(OH)2) = n(OH-)/2 = 0.4 mol / 2 = 0.2 mol


Calculate the mass of Cu(OH)2:

The molar mass of Cu(OH)2 is 97.56 g/mol.

Therefore,

(0.2 mol Cu(OH)2) × (97.56 g Cu(OH)2 / 1 mol Cu(OH)2) = 19.512 g Cu(OH)2 = 19.5 g Cu(OH)2

Mass of Cu(OH)2 is 19.5 g.


Answer: 19.5 grams of copper(II) hydroxide (Cu(OH)2) would be produced.